OpenStudy (anonymous):

Potassium-40 has a half-life of approximately 1.25 billion years. Approximately how many years will pass before a sample of potassium-40 contains one-eighth the original amount of parent isotope?

OpenStudy (frostbite):

You know the radioactive decay law?

OpenStudy (anonymous):

No, please explain.

OpenStudy (frostbite):

Right. Well isotopes decay due to a exponential function. The differential equation that should be solved is: \[\LARGE -\frac{ dN }{ N }= \lambda dt\] The differential question has the solution: \[\LARGE N(t)=N _{0}e ^{-\lambda*t}\] Where the half-life is given by: \[\LARGE t _{1/2}=\frac{ \ln(2) }{\lambda }\]

OpenStudy (frostbite):

Now lets set up the problem: we want to know how much time there has went where there are 1/8 of the sample left. So the problem most be: \[\LARGE N(t)=\frac{ 1 }{ 8 }N _{0}\] The radioactive decay law then become: \[\LARGE \frac{ 1 }{ 8 }N _{0}=N_{0}e ^{-\lambda*t}\] The N_0 cancel on both sides: \[\LARGE \frac{ 1 }{ 8 }=e ^{-\lambda*t}\] We substitute the expression for the half-life: \[\LARGE \frac{ 1 }{ 8 }=e ^{\frac{ \ln(2) }{ t _{1/2}}*t}\] Solve for t and you have the answer.

OpenStudy (anonymous):

@Yairenis Do you know what "half life" means?

OpenStudy (frostbite):

Woops a minus sign is missing in the last expression.

OpenStudy (frostbite):

Anyway think I'm gonna solve for you, just to make this one an example for future work: \[\LARGE \frac{ 1 }{ 8 }=e ^{-\frac{ \ln(2) }{ t _{1/2} }*t}\] \[\LARGE \ln \left( \frac{ 1 }{ 8 } \right)=-\frac{ \ln(2) }{ t _{1/2} }*t\] \[\LARGE t=-\frac{ \ln \left( \frac{ 1 }{ 8 } \right) }{ \frac{ \ln \left( 2 \right) }{ t _{1/2} } }\]

OpenStudy (frostbite):

Don't worry about the minus sign... the natural logarithm to a number below 1 is negative, so the time is going to get positive.

OpenStudy (anonymous):

So the answer would be 3.75 billion?

OpenStudy (anonymous):

Yes please

OpenStudy (frostbite):

Yea I get the same.

OpenStudy (anonymous):

Then thank you for your help, Frostbite

OpenStudy (frostbite):

No problem at all. Hope the outline was understandable?

OpenStudy (anonymous):

Yes

OpenStudy (frostbite):

Perfect.